OpenStudy (anonymous):

(x -5)2 + (y + 2)2 = 64 The radius of the circle is ________ units.

OpenStudy (anonymous):

whats the general equation of a circle?

OpenStudy (anonymous):

(x-h)^2+(y-k)^2 = r^2 what is the radius ?

OpenStudy (anonymous):

@surdawi has given you the answer to what I asked.

OpenStudy (anonymous):

x^2 - 25 + y^2 + 4 = 64 ... am i do this right so far? @Brittni0605 & @surdawi

OpenStudy (anonymous):

You're going to far. YOU just want the radius

OpenStudy (anonymous):

you dont have to distribute (x-h)^2+(y-k)^2 = r^2 (x -5)2 + (y + 2)2 = 64 what is the radius

OpenStudy (anonymous):

so its 64

OpenStudy (anonymous):